OpenStudy (anonymous):

Verify the Identity cos(α - β) - cos(α + β) = 2 sin α sin β @Nnesha

OpenStudy (anonymous):

cos=1/sec

OpenStudy (anonymous):

1/sec(a-b)-1/sec(a+b)=2sin(a) sin(b)

OpenStudy (anonymous):

i cant read what you wrote lol

OpenStudy (anonymous):

why do i have a feeling everything iv learned about verifying so far will not help me here

OpenStudy (butterflydreamer):

cos(α - β) = cosα cosβ + sinα sin β cos(α + β) = cosα cosβ - sinα sin β These are two rules which you must know :) They're known as your addition and subtraction (or sum and difference) formulas for cosine

OpenStudy (anonymous):

uhhhh i see a bunch of question marks on my screen^

rvc (rvc):

trigo is important

OpenStudy (butterflydreamer):

refresh your page

rvc (rvc):

wait\[\cos(a+b)=cosacosb-sinasinb\] \[\cos(a-b)=cosacosb+sinasinb\]

OpenStudy (butterflydreamer):

yyaasss :D So those are the two formulas you need to remember @quick24

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can yall give me like 3 minutes i need a smoke break

rvc (rvc):

for cos its opposite signs whereas for sine its the same sign t\[\sin(a+b)=sinacosb+cosasinb\] \[\sin(a-b)=sinacosb-cosasinb\]

OpenStudy (anonymous):

ok im backn @rvc @butterflydreamer

OpenStudy (anonymous):

hey rvc

rvc (rvc):

so u got those formulas?

OpenStudy (anonymous):

i havent copied them yet but i will

rvc (rvc):

do it pls

OpenStudy (anonymous):

i copied cos(a+b)=cosacosb−sinasinb and cos(a−b)=cosacosb+sinasinb r those the ones i need

OpenStudy (anonymous):

@rvc

rvc (rvc):

yes

OpenStudy (anonymous):

so now what

rvc (rvc):

what u need?

OpenStudy (anonymous):

i need to verify L.H.S=R.H.S

OpenStudy (anonymous):

can you walk me through it and yes i will do the lifting just need you to give me hints in the right direction

rvc (rvc):

yes what help u need

OpenStudy (anonymous):

just start from the begging and explain what i have to do like how to apply those formulas for instance

rvc (rvc):

well what grade r u in?

OpenStudy (anonymous):

12th

rvc (rvc):

so that i may direct u in the right way

OpenStudy (anonymous):

im a senior

OpenStudy (anonymous):

no

OpenStudy (anonymous):

why

rvc (rvc):

okay just to get a idea what kinda questions u may get

OpenStudy (anonymous):

im white in america and im a senior

rvc (rvc):

hmm.. look to start with u may gett this to prove lhs or rhs

OpenStudy (anonymous):

im to prove rhs im assuming as thats what i had to do in the other verification problem

rvc (rvc):

oh

rvc (rvc):

expand the LHS @quick24

OpenStudy (anonymous):

so plug in that equation to replace the L.H.S terms

rvc (rvc):

yes!

rvc (rvc):

show the working @quick24

OpenStudy (anonymous):

ok so its (cos(a)*cos(b)+sin(a)*sin(b))-(cos(a)*cos(b)-sin(a)*sin(b))

OpenStudy (anonymous):

[(\cos(a)*\cos(b)+\sin(a)*\sin(b)) - (\cos(a)*\cos(b)-\sin(a)*\sin(b)\]

OpenStudy (anonymous):

\[(\cos(a)*\cos(b)+\sin(a)*\sin(b))-(\cos(a)*\cos(b)-\sin(a)*\sin(b))\]

rvc (rvc):

hey since its has angles alpha and beta use them instead of a nd b

rvc (rvc):

now open the brackets

OpenStudy (anonymous):

will you still be on in about 30 minutes i have to drive my grandmother somewhere

OpenStudy (anonymous):

rvc?

rvc (rvc):

okay

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

@rvc hey